On the Standford-Binet test, the mean IQ is 100. A class of 21 kindergarten pupils were tested with a resulting mean of 108 and a standard deviation of 7.07. What is the upper bound of the 90% confidence interval of the average IQ of the kindergarten pupils?



1
Expert's answer
2022-03-01T12:51:51-0500

Upper bound: "\bar x+t_{0.05,n-1}\frac{s}{\sqrt{n}}=108+1.724\frac{7.07}{\sqrt{21}}=110.66."